# Optimal frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v1

@@ 3.1. The Relationship between Tight Frames and Convex Polygons in $\mathbb{R}^{2}$
Figure 3. One of the optimal dodecagon as in Figure 2 (a2), with vertices $\bm{v}_{1},\dotsc,\bm{v}_{12}$. The corresponding Reuleaux polygon consists of $9$ circular arcs with vertices $\bm{v}_{c_{1}},\dotsc,\bm{v}_{c_{9}}$ indicated in red dots. The dashed lines connect pairs of vertices at maximal distance.
@@
Assume that $m\geq 3$ is an integer with an odd factor. Let $P$ be a convex $m$-gon with edge vectors $\{\bm{e}_{1},\dotsc\bm{e}_{m}\}$. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from $\bm{e}_{1}$ to $\bm{e}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. Let ${\mathcal{E}}_{m}$ denote the collection of all optimal strictly convex $m$-gons in $\mathbb{R}^{2}$. If $P\in{\mathcal{E}}_{m}$ then
@@
Assume that $P\in{\mathcal{E}}_{m}$. It follows from Theorem 4.2 that $P$ is equilateral. We now prove that $P$ satisfies (4.1). Assume without loss of generality that $\bm{e}_{1},\dotsc,\bm{e}_{m}$ are arranged counterclockwise. Let $\bm{v}_{1},\dotsc,\bm{v}_{m}$ be the vertices of $P$ such that $\bm{e}_{j}=\bm{v}_{j+1}-\bm{v}_{j}$ for each $j\in[m-1]$ and $\bm{e}_{m}=\bm{v}_{1}-\bm{v}_{m}$. For each $i\in[m]$, we denote $\bm{v}_{m+i}=\bm{v}_{i}$, and let $\theta_{i}$ be the counterclockwise angle from $\bm{e}_{i-1}$ to $\bm{e}_{i}$, where $\bm{e}_{0}:=\bm{e}_{m}$. Then we have $\alpha_{1}=0$ and $\alpha_{i}=\alpha_{i-1}+\theta_{i}$ for each $i\in\{2,\ldots,m\}$. Moreover, since $P$ is strictly convex, we have $\theta_{i}\in(0,\pi)$, $\forall i\in[m]$.
@@
By Theorem 4.1 (ii), for each $i\in[r]$, $\bm{v}_{c_{i}}$ is the center of the circular arc ¿ $\bm{v}_{c_{i+h}}\bm{v}_{c_{i+h+1}}$ , where $h=\frac{r-1}{2}$ is a fixed integer. See Figure 3 for a clear and intuitive illustration. Since $P$ is equilateral and the sum of the interior angles at the vertices of $S$ equals $\pi$, the interior angle of the arc ¿ $\bm{v}_{c_{i+h}}\bm{v}_{c_{i+h+1}}$ centered on $\bm{v}_{c_{i}}$ is $(c_{i+h+1}-c_{i+h})\cdot\frac{\pi}{m}$ for each $i\in[r]$. By plane geometry, if $i\notin\{c_{1},\ldots,c_{r}\}$ then we have
@@
(i) We first prove that there exists a convex $m$-gon $P$ with the edge set $\{\varepsilon_{1}\bm{\mu}_{1},\dotsc,\varepsilon_{m}\bm{\mu}_{m}\}$. Since $\bm{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}$ satisfies (1.11), we have
@@
$\sum_{j=1}^{m}\varepsilon_{j}\bm{\mu}_{j}=(\sum_{j=1}^{m}\varepsilon_{j}\cos\frac{j\pi}{m},\sum_{j=1}^{m}\varepsilon_{j}\sin\frac{j\pi}{m})^{T}=\bm{0}.$ (4.9)
Combining with the fact that $\varepsilon_{1}\bm{\mu}_{1},\ldots,\varepsilon_{m}\bm{\mu}_{m}$ have pairwise distinct directions, we can utilize Corollary 2.1 to obtain a strictly convex $m$-gon $P$ with the edge set $\{\varepsilon_{1}\bm{\mu}_{1},\dotsc,\varepsilon_{m}\bm{\mu}_{m}\}$, and $P$ is unique up to translation.
We next prove that $P\in{\mathcal{E}}_{m}$. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from $\varepsilon_{1}\bm{\mu}_{1}$ to $\varepsilon_{i}\bm{\mu}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. Since each $\bm{\mu}_{j}=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$, we have
@@
Let $Q\in{\mathcal{E}}_{m}$ be an optimal $m$-polygon with the edge set $\{\bm{q}_{1},\dotsc,\bm{q}_{m}\}$ such that $\bm{q}_{1}=\varepsilon_{1}\bm{\mu}_{1}$. By Theorem 4.2 we see that $Q$ is equilateral, so we have $\|\bm{q}_{j}\|_{2}=1$ for each $j\in[m]$. For each $i\in[m]$, define $\beta_{i}\in[0,2\pi)$ as the counterclockwise angle from $\bm{q}_{1}$ to $\bm{q}_{i}$, and let $\phi_{i}\in(0,\pi]$ satisfy $\phi_{i}\equiv\beta_{i}\pmod{\pi}$. By Lemma 4.1 we have
@@
Therefore, there exists a permutation $\{s_{1},s_{2},\ldots,s_{m}\}$ of $[m]$ such that $\psi_{i}=\phi_{s_{i}}$ for each $i\in[m]$. This means that for each $i\in[m]$ we have $\alpha_{i}\equiv\beta_{s_{i}}\pmod{\pi}$. Recall that $\bm{q}_{1}=\varepsilon_{1}\bm{\mu}_{1}=\varepsilon_{1}(\cos\frac{\pi}{m},\sin\frac{\pi}{m})^{T}$. Let $I:=\{i\in[m]\mid\alpha_{i}=\beta_{s_{i}}\}$. If $i\in I$, then
$\bm{q}_{s_{i}}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\beta_{s_{i}}),\sin(\frac{\pi}{m}+\beta_{s_{i}}))^{T}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\alpha_{i}),\sin(\frac{\pi}{m}+\alpha_{i}))^{T}=\varepsilon_{i}\bm{\mu}_{i}.$
Otherwise, if i[$\Ii\in[m]\backslash$ I, then $\alpha_{i}=\beta_{s_{i}}-\pi$ or $\alpha_{i}=\beta_{s_{i}}+\pi$. So we have
$\bm{q}_{s_{i}}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\beta_{s_{i}}),\sin(\frac{\pi}{m}+\beta_{s_{i}}))^{T}=-\varepsilon_{1}(\cos(\frac{\pi}{m}+\alpha_{i}),\sin(\frac{\pi}{m}+\alpha_{i}))^{T}=-\varepsilon_{i}\bm{\mu}_{i}.$
Hence, we have $\{\bm{q}_{1},\dotsc,\bm{q}_{m}\}=\{\varepsilon_{i}\bm{\mu}_{i}\}_{i\in I}\cup\{-\varepsilon_{i}\bm{\mu}_{i}\}_{i\in[m]\setminus I}$. According to (2.7) in Lemma 2.4, we have
$r({P})=\frac{\max_{\bm{u}\in\mathbb{S}^{1}}\,\sum_{i=1}^{m}{|\langle\bm{\mu}_{i},\bm{u}\rangle|}}{2\sum_{i=1}^{m}\|\bm{\mu}_{i}\|_{2}}=\frac{\max_{\bm{u}\in\mathbb{S}^{1}}\,\sum_{i=1}^{m}{|\langle\bm{q}_{i},\bm{u}\rangle|}}{2\sum_{i=1}^{m}\|\bm{q}_{i}\|_{2}}=r(Q).$
@@
(ii) Let $P\in{\mathcal{E}}_{m}$ be an optimal polygon with edge vectors $\{{\bm{e}}_{1},\dotsc,{\bm{e}}_{m}\}$. Without loss of generality, we may assume that ${\bm{e}}_{1}=(1,0)^{T}$ after a proper scaling and rotation. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from ${\bm{e}}_{1}$ to ${\bm{e}}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. By Lemma 4.1, we see that $P$ is equilateral and
@@
Hence, there exists a permutation $\{s_{1},\ldots,s_{m}\}$ of $[m]$ such that $\psi_{s_{j}}=\frac{j\pi}{m}$ for each $j\in[m]$. Let $I:=\{j\in[m]:\alpha_{s_{j}}=\psi_{s_{j}}\}$, and let $\bm{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}$, where $\varepsilon_{j}=1$ if $j\in I$ and $\varepsilon_{j}=-1$ if $j\notin I$. Recall that ${\bm{e}}_{1}=(1,0)^{T}$ and $\alpha_{i}$ is the counterclockwise angle from ${\bm{e}}_{1}$ to ${\bm{e}}_{i}$. Then we have
${\bm{e}}_{s_{j}}$ $=(\cos\alpha_{s_{j}},\sin\alpha_{s_{j}})^{T}=(\cos\psi_{s_{j}},\sin\psi_{s_{j}})^{T}=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}=\bm{\mu}_{j}\quad\text{for}\;j\in I,$
${\bm{e}}_{s_{j}}$ $=(\cos\alpha_{s_{j}},\sin\alpha_{s_{j}})^{T}=(\cos(\pi+\psi_{s_{j}}),\sin(\pi+\psi_{s_{j}}))^{T}=-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}=-\bm{\mu}_{j}\quad\text{for}\;j\in[m]\setminus I.$
Therefore, the edge vectors of $P$ are $\{\varepsilon_{1}\bm{\mu}_{1},\ldots,\varepsilon_{m}\bm{\mu}_{m}\}$.
It suffices to prove that $\bm{\varepsilon}$ satisfies (1.11). By Corollary 2.1, we have
$\bm{0}=\sum_{i=1}^{m}\varepsilon_{i}\bm{\mu}_{i}=(\sum_{j=1}^{m}\varepsilon_{j}\cos\frac{j\pi}{m},\sum_{j=1}^{m}\varepsilon_{j}\sin\frac{j\pi}{m})^{T},$
@@
[32] Let $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$. Then
$(L_{\bm{A}})^{2}\leq\frac{1}{2}\sum_{j=1}^{m}\|\bm{a}_{j}\|_{2}^{2}-\frac{1}{2m\cdot\sin\frac{\pi}{2m}}\sum_{j=1}^{m}\|\bm{a}_{j}\|_{2}^{2},$
where $L_{\bm{A}}$ is the lower Lipschitz constant of the map $\Phi_{\bm{A}}$ as defined in (1.1).
@@
(i) We first prove that $\mathcal{A}$ forms a tight frame by contradiction. Suppose, for contradiction, that $\mathcal{A}$ is not a tight frame. Set $\bm{A}=[\bm{a}_{1},\ldots,\bm{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$. By Theorem 2.2, we have
$(U_{\bm{A}})^{2}=\|\bm{A}\|_{2}^{2}=\|\bm{A}^{T}\bm{A}\|_{2}>\frac{1}{2}\cdot\operatorname{Tr}(\bm{A}^{T}\bm{A})=\frac{1}{2}\cdot\operatorname{Tr}(\bm{A}\bm{A}^{T})=\frac{1}{2}\sum_{j=1}^{m}\|\bm{a}_{j}\|_{2}^{2}.$
@@
$\beta_{\bm{A}}=\frac{U_{\bm{A}}}{L_{\bm{A}}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (4.12)
@@
Corollary 1.1 (i) shows that $\min_{\bm{M}\in\mathbb{R}^{m\times 2}}\beta_{\bm{M}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}$. Combining with (4.12), we obtain that $\beta_{\bm{A}}>\min_{\bm{M}\in\mathbb{R}^{m\times 2}}\beta_{\bm{M}}$, meaning that $\mathcal{A}$ is not optimal. This contradicts with $\mathcal{A}\in\mathcal{T}_{m}$. Therefore, $\mathcal{A}$ forms a tight frame.
@@
We next prove that all vectors in $\mathcal{A}$} have the same norm. By Theorem 1.2, there is an optimal strictly convex $m$-gon $P$ with the edge set $\{f(\bm{a}_{1}),\dotsc,f(\bm{a}_{m})\}$. Since $m\geq 3$ is an integer with an odd factor, Theorem 1.1 (i) implies that $P$ is equilateral, i.e., $\|f(\bm{a}_{1})\|_{2}=\|f(\bm{a}_{2})\|_{2}=\cdots=\|f(\bm{a}_{m})\|_{2}$. Consequently, according to the definition of $f(\cdot)$ in (1.10), all vectors in $\mathcal{A}$ have equal norm.

### v2

@@ 3.1. The relationship between tight frames and convex polygons in $\mathbb{R}^{2}$
Remark 4.2.
Let $P$ be an optimal convex $m$-gon, where $m\geq 3$ is an integer with an odd factor. By Theorem 4.2 and Theorem 4.1 (i), $P$ is inscribed in a Reuleaux polygon $R$ with $r$ boundary arcs, where $r$ is an odd integer with $3\leq r\leq m$. Label the vertices of $R$ as $\boldsymbol{w}_{1},\dotsc,\boldsymbol{w}_{r}$ in counterclockwise order, and extend the indexing cyclically by setting $\boldsymbol{w}_{r+i}=\boldsymbol{w}_{i}$ for each $i\in[r]$. Define $h=\frac{r-1}{2}$. Then for each $i\in[r]$, the vertex $\boldsymbol{w}_{i}$ is the the center of the arc ¿ $\boldsymbol{w}_{{i+h}}\boldsymbol{w}_{{i+h+1}}$ on the boundary of $R$. This can be seen as follows. By Theorem 4.1 (ii), we can start at $\boldsymbol{s}_{1}:=\boldsymbol{w}_{1}$ and construct a star polygon $S$ by sequentially connecting pairs of vertices in $R$ at maximal distance in the counterclockwise direction. Let $\boldsymbol{s}_{i}$ denote the $i$-th vertex of $R$ visited during this construction. According to [25, Page 256], the vertices $\boldsymbol{s}_{1},\boldsymbol{s}_{3},\dotsc,\boldsymbol{s}_{r},\boldsymbol{s}_{2},\boldsymbol{s}_{4},\dotsc,\boldsymbol{s}_{r-1}$ are ordered counterclockwise. Comparing this with the labeling $\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{r}$, we have
$\boldsymbol{w}_{1}=\boldsymbol{s}_{1},\boldsymbol{w}_{2}=\boldsymbol{s}_{3},\ldots,\boldsymbol{w}_{\frac{r+1}{2}}=\boldsymbol{s}_{r},\boldsymbol{w}_{\frac{r+1}{2}+1}=\boldsymbol{s}_{2},\ldots,\boldsymbol{w}_{r}=\boldsymbol{s}_{r-1}.$
Set $\boldsymbol{s}_{0}:=\boldsymbol{s}_{r}$ and $\boldsymbol{s}_{r+1}:=\boldsymbol{s}_{1}$. Note that $\boldsymbol{s}_{i}$ is the center of the arc ¿ $\boldsymbol{s}_{i-1}\boldsymbol{s}_{i+1}$ for each $i\in[r]$. Translating this to the $\boldsymbol{w}_{i}$ notation, we conclude that for each $i\in[r]$, $\boldsymbol{w}_{{i}}$ is the center of the arc ¿ $\boldsymbol{w}_{{i+h}}\boldsymbol{w}_{{i+h+1}}$ , where $h=\frac{r-1}{2}$.
Figure 3. One of the optimal dodecagon as in Figure 2 (a2), with vertices $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{12}$. The corresponding Reuleaux polygon consists of $9$ circular arcs with vertices $\boldsymbol{v}_{c_{1}},\dotsc,\boldsymbol{v}_{c_{9}}$ indicated in red dots. The dashed lines connect pairs of vertices at maximal distance.
@@
Assume that $m\geq 3$ is an integer with an odd factor. Let $P$ be a convex $m$-gon with edge vectors $\{\boldsymbol{e}_{1},\dotsc\boldsymbol{e}_{m}\}$. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from $\boldsymbol{e}_{1}$ to $\boldsymbol{e}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. Let ${\mathcal{E}}_{m}$ denote the collection of all optimal strictly convex $m$-gons in $\mathbb{R}^{2}$. If $P\in{\mathcal{E}}_{m}$ then
@@
Assume that $P\in{\mathcal{E}}_{m}$. It follows from Theorem 4.2 that $P$ is equilateral. We now prove that $P$ satisfies (4.1). Assume without loss of generality that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. Let $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ be the vertices of $P$ such that $\boldsymbol{e}_{j}=\boldsymbol{v}_{j+1}-\boldsymbol{v}_{j}$ for each $j\in[m-1]$ and $\boldsymbol{e}_{m}=\boldsymbol{v}_{1}-\boldsymbol{v}_{m}$. For each $i\in[m]$, we denote $\boldsymbol{v}_{m+i}=\boldsymbol{v}_{i}$, and let $\theta_{i}$ be the counterclockwise angle from $\boldsymbol{e}_{i-1}$ to $\boldsymbol{e}_{i}$, where $\boldsymbol{e}_{0}:=\boldsymbol{e}_{m}$. Then we have $\alpha_{1}=0$ and $\alpha_{i}=\alpha_{i-1}+\theta_{i}$ for each $i\in\{2,\ldots,m\}$. Moreover, since $P$ is strictly convex, we have $\theta_{i}\in(0,\pi)$, $\forall i\in[m]$.
@@
By Remark 4.2, for each $i\in[r]$, $\boldsymbol{v}_{c_{i}}$ is the center of the arc ¿ $\boldsymbol{v}_{c_{i+h}}\boldsymbol{v}_{c_{i+h+1}}$ , where $h=\frac{r-1}{2}$ is a fixed integer. See Figure 3 for a clear and intuitive illustration. Since $P$ is equilateral and the sum of the interior angles at the vertices of $R$ equals $\pi$, the interior angle of the arc ¿ $\boldsymbol{v}_{c_{i+h}}\boldsymbol{v}_{c_{i+h+1}}$ centered on $\boldsymbol{v}_{c_{i}}$ is $(c_{i+h+1}-c_{i+h})\cdot\frac{\pi}{m}$ for each $i\in[r]$. By plane geometry, if $i\notin\{c_{1},\ldots,c_{r}\}$ then we have
@@
(i) We first prove that there exists a convex $m$-gon $P$ with the edge set $\{\varepsilon_{1}\boldsymbol{\mu}_{1},\dotsc,\varepsilon_{m}\boldsymbol{\mu}_{m}\}$. Since $\boldsymbol{\varepsilon}=(\varepsilon_{1},\ldots,\varepsilon_{m})^{T}$ satisfies (1.11), we have
@@
$\sum_{j=1}^{m}\varepsilon_{j}\boldsymbol{\mu}_{j}=(\sum_{j=1}^{m}\varepsilon_{j}\cos\frac{j\pi}{m},\sum_{j=1}^{m}\varepsilon_{j}\sin\frac{j\pi}{m})^{T}=\boldsymbol{0}.$ (4.9)
Combining with the fact that $\varepsilon_{1}\boldsymbol{\mu}_{1},\ldots,\varepsilon_{m}\boldsymbol{\mu}_{m}$ have pairwise distinct directions, we can utilize Corollary 2.1 to obtain a strictly convex $m$-gon $P$ with the edge set $\{\varepsilon_{1}\boldsymbol{\mu}_{1},\dotsc,\varepsilon_{m}\boldsymbol{\mu}_{m}\}$, and $P$ is unique up to translation.
We next prove that $P\in{\mathcal{E}}_{m}$. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from $\varepsilon_{1}\boldsymbol{\mu}_{1}$ to $\varepsilon_{i}\boldsymbol{\mu}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. Since each $\boldsymbol{\mu}_{j}=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$, we have
@@
Let $Q\in{\mathcal{E}}_{m}$ be an optimal $m$-polygon with the edge set $\{\boldsymbol{q}_{1},\dotsc,\boldsymbol{q}_{m}\}$ such that $\boldsymbol{q}_{1}=\varepsilon_{1}\boldsymbol{\mu}_{1}$. By Theorem 4.2 we see that $Q$ is equilateral, so we have $\|\boldsymbol{q}_{j}\|_{2}=1$ for each $j\in[m]$. For each $i\in[m]$, define $\beta_{i}\in[0,2\pi)$ as the counterclockwise angle from $\boldsymbol{q}_{1}$ to $\boldsymbol{q}_{i}$, and let $\phi_{i}\in(0,\pi]$ satisfy $\phi_{i}\equiv\beta_{i}\pmod{\pi}$. By Lemma 4.1 we have
@@
Therefore, there exists a permutation $\{s_{1},s_{2},\ldots,s_{m}\}$ of $[m]$ such that $\psi_{i}=\phi_{s_{i}}$ for each $i\in[m]$. This means that for each $i\in[m]$ we have $\alpha_{i}\equiv\beta_{s_{i}}\pmod{\pi}$. Recall that $\boldsymbol{q}_{1}=\varepsilon_{1}\boldsymbol{\mu}_{1}=\varepsilon_{1}(\cos\frac{\pi}{m},\sin\frac{\pi}{m})^{T}$. Let $I:=\{i\in[m]\mid\alpha_{i}=\beta_{s_{i}}\}$. If $i\in I$, then
$\boldsymbol{q}_{s_{i}}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\beta_{s_{i}}),\sin(\frac{\pi}{m}+\beta_{s_{i}}))^{T}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\alpha_{i}),\sin(\frac{\pi}{m}+\alpha_{i}))^{T}=\varepsilon_{i}\boldsymbol{\mu}_{i}.$
Otherwise, if $i\in[m]\setminus I$, then $\alpha_{i}=\beta_{s_{i}}-\pi$ or $\alpha_{i}=\beta_{s_{i}}+\pi$. So we have
$\boldsymbol{q}_{s_{i}}=\varepsilon_{1}(\cos(\frac{\pi}{m}+\beta_{s_{i}}),\sin(\frac{\pi}{m}+\beta_{s_{i}}))^{T}=-\varepsilon_{1}(\cos(\frac{\pi}{m}+\alpha_{i}),\sin(\frac{\pi}{m}+\alpha_{i}))^{T}=-\varepsilon_{i}\boldsymbol{\mu}_{i}.$
Hence, we have $\{\boldsymbol{q}_{1},\dotsc,\boldsymbol{q}_{m}\}=\{\varepsilon_{i}\boldsymbol{\mu}_{i}\}_{i\in I}\cup\{-\varepsilon_{i}\boldsymbol{\mu}_{i}\}_{i\in[m]\setminus I}$. According to (2.7) in Lemma 2.4, we have
$r({P})=\frac{\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,\sum_{i=1}^{m}{|\langle\boldsymbol{\mu}_{i},\boldsymbol{u}\rangle|}}{2\sum_{i=1}^{m}\|\boldsymbol{\mu}_{i}\|_{2}}=\frac{\max_{\boldsymbol{u}\in\mathbb{S}^{1}}\,\sum_{i=1}^{m}{|\langle\boldsymbol{q}_{i},\boldsymbol{u}\rangle|}}{2\sum_{i=1}^{m}\|\boldsymbol{q}_{i}\|_{2}}=r(Q).$
@@
(ii) Let $P\in{\mathcal{E}}_{m}$ be an optimal polygon with edge vectors $\{{\boldsymbol{e}}_{1},\dotsc,{\boldsymbol{e}}_{m}\}$. Without loss of generality, we may assume that ${\boldsymbol{e}}_{1}=(1,0)^{T}$ after a proper scaling and rotation. For each $i\in[m]$, define $\alpha_{i}\in[0,2\pi)$ as the counterclockwise angle from ${\boldsymbol{e}}_{1}$ to ${\boldsymbol{e}}_{i}$, and let $\psi_{i}\in(0,\pi]$ satisfy $\psi_{i}\equiv\alpha_{i}\pmod{\pi}$. By Lemma 4.1, we see that $P$ is equilateral and
@@
Hence, there exists a permutation $\{s_{1},\ldots,s_{m}\}$ of $[m]$ such that $\psi_{s_{j}}=\frac{j\pi}{m}$ for each $j\in[m]$. Let $I:=\{j\in[m]:\alpha_{s_{j}}=\psi_{s_{j}}\}$, and let $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}$, where $\varepsilon_{j}=1$ if $j\in I$ and $\varepsilon_{j}=-1$ if $j\notin I$. Recall that ${\boldsymbol{e}}_{1}=(1,0)^{T}$ and $\alpha_{i}$ is the counterclockwise angle from ${\boldsymbol{e}}_{1}$ to ${\boldsymbol{e}}_{i}$. Then we have
${\boldsymbol{e}}_{s_{j}}=(\cos\alpha_{s_{j}},\sin\alpha_{s_{j}})^{T}$ $=(\cos\psi_{s_{j}},\sin\psi_{s_{j}})^{T}=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}=\boldsymbol{\mu}_{j}\quad\text{for}\;j\in I,$
${\boldsymbol{e}}_{s_{j}}=(\cos\alpha_{s_{j}},\sin\alpha_{s_{j}})^{T}$ $=(\cos(\pi+\psi_{s_{j}}),\sin(\pi+\psi_{s_{j}}))^{T}$
$=-(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}=-\boldsymbol{\mu}_{j}\quad\text{for}\;j\in[m]\setminus I.$
Therefore, the edge vectors of $P$ are $\{\varepsilon_{1}\boldsymbol{\mu}_{1},\ldots,\varepsilon_{m}\boldsymbol{\mu}_{m}\}$.
It suffices to prove that $\boldsymbol{\varepsilon}$ satisfies (1.11). By Corollary 2.1, we have
$\boldsymbol{0}=\sum_{i=1}^{m}\varepsilon_{i}\boldsymbol{\mu}_{i}=(\sum_{j=1}^{m}\varepsilon_{j}\cos\frac{j\pi}{m},\sum_{j=1}^{m}\varepsilon_{j}\sin\frac{j\pi}{m})^{T},$
@@
[32] Let $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}\in\mathbb{R}^{m\times 2}$. Then
$(L_{\boldsymbol{A}})^{2}\leq\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2}-\frac{1}{2m\cdot\sin\frac{\pi}{2m}}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2},$
where $L_{\boldsymbol{A}}$ is the optimal lower Lipschitz constant of the map $\Phi_{\boldsymbol{A}}$ defined in (1.1).
@@
(i) We first prove that $\mathcal{A}$ forms a tight frame by contradiction. Suppose, for contradiction, that $\mathcal{A}$ is not a tight frame. Set $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}$ with $\boldsymbol{A}\in\mathbb{R}^{m\times 2}$. By Theorem 2.2, we have
$(U_{\boldsymbol{A}})^{2}=\|\boldsymbol{A}\|_{2}^{2}=\|\boldsymbol{A}^{T}\boldsymbol{A}\|_{2}>\frac{1}{2}\cdot\operatorname{Tr}(\boldsymbol{A}^{T}\boldsymbol{A})=\frac{1}{2}\cdot\operatorname{Tr}(\boldsymbol{A}\boldsymbol{A}^{T})=\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2}.$
@@
$\beta_{\boldsymbol{A}}=\frac{U_{\boldsymbol{A}}}{L_{\boldsymbol{A}}}>\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}.$ (4.12)
@@
Corollary 1.1 (i) shows that $\min_{\boldsymbol{M}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{M}}=\frac{1}{\sqrt{1-\frac{1}{m\cdot\sin\frac{\pi}{2m}}}}$. Combining with (4.12), we obtain that $\beta_{\boldsymbol{A}}>\min_{\boldsymbol{M}\in\mathbb{R}^{m\times 2}}\beta_{\boldsymbol{M}}$, meaning that $\mathcal{A}$ is not optimal. This contradicts with $\mathcal{A}\in\mathcal{T}_{m}$. Therefore, $\mathcal{A}$ forms a tight frame.
@@
We next prove that all vectors in $\mathcal{A}$} have the same norm. By Theorem 1.2, there is an optimal strictly convex $m$-gon $P$ with the edge set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$. Since $m\geq 3$ is an integer with an odd factor, Theorem 1.1 (i) implies that $P$ is equilateral, i.e., $\|f(\boldsymbol{a}_{1})\|_{2}=\|f(\boldsymbol{a}_{2})\|_{2}=\cdots=\|f(\boldsymbol{a}_{m})\|_{2}$. Consequently, according to the definition of $f(\cdot)$ in (1.10), all vectors in $\mathcal{A}$ have equal norm.